A plane flies 420 miles with the wind and 300 miles against the wind in the same length of time. If the speed of the wind is 23 ​mph, what is the speed of the plane in still​ air?
Answer by stanbon(75887)

A plane flies 420 miles with the wind and 300 miles against the wind in the same length of time. If the speed of the wind is 23 ​mph, what is the speed of the plane in still​ air?
--------------------
With-wind DATA ::
dist = 420 miles ; rate = p+23 ; time = d/r = 420/(p+23) hrs
-----------------
Against-wind DATA:
dist = 300 miles ; rate = p-23 ; time = d/r = 300/(p-23) hrs
---------------------
Equation:
420/(p+23) = 300/(p-23)
---------
7/(p+23) = 5/(p-23)
--------------------------
7p-161 = 5p+115
----------------------
2p = 276
------
p = 138 mph (speed of the plane in still air)
------------
Cheers,
Stan H.
----------------